OpenStudy (axel.caballero):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4 (1 point) 2(5y - 1) + 4y = -4 2(-5y + 1) + 4y = -4 2x + 4(-5y + 1) = -4 2x + 4(5y - 1) = -4

OpenStudy (axel.caballero):

Can anyone help ? @satellite73 @Hero @ganeshie8 @TuringTest @Nurali @ash2326 @Luigi0210 @abb0t @kelliegirl33 @eliassaab @ranga @blues @ashleyisakitty @nubeer @FutureMathProfessor @shrutipande9 @chooch146 @allie_bear22

OpenStudy (nurali):

-x + 5y = 1----(1) 2x + 4y = -4---(2) In Equation (1) -x + 5y = 1 x=5y-1 put 'x' in equation (2) 2x + 4y = -4 2(5y-1)+4y=-4